OpenStudy (anonymous):

$80.00 is 25% of what number? please explain how to get the answer.

OpenStudy (anonymous):

320 .25x = 80 x= 80 /.25 = 320

OpenStudy (anonymous):

255 = 1/4, so 80 is 1/4 of the number so multiply it by 4.

OpenStudy (anonymous):

320. you set up a proportion. % means per 100. 80/x=25/100. you cross multiply then divide so 100 times 80 which is 8000 then divide that by 25. which equals 320

OpenStudy (anonymous):

thank you soooo much, i was having trouble on that question.